1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?
Answer by toni16951(1)

x+8(2)=56
x+ 16 =56
-16 =-16
x=40
So that means that Belinda is 40 years old. You know that the two combined ages have to add up tp 56 years so 56-40=16 which means that Tony is 16 years of age. To double check your answer you can start with Tonys age and it states that Belinda is twice Tony's age plus 16 so you can take Tony's age 16 multipy it by 2 which gives you 32 and then you add 8 which gives you 40. So fourty and 16 equals 56 and when we double check it (16x2+8=40) both equations work for this problem.
Belinda=40 years of age
Tony=16 years of age